Diamond and Turf Inc. is considering an investment in one of two machines. The sewing machine will increase productivity from sewing 140 baseballs per hour to sewing 252 per hour. The contribution margin per unit is $0.42 per baseball. Assume that any increased production of baseballs can be sold. The second machine is an automatic packing machine for the golf ball line. The packing machine will reduce packing labor cost. The labor cost saved is equivalent to $26 per hour. The sewing machine will cost $266,100, will have a seven-year life, and will operate for 1,400 hours per year. The packing machine will cost $132,100, will have a seven-year life, and will operate for 1,200 hours per year. Diamond and Turf seeks a minimum rate of return of 10% on its investments.

Present Value of an Annuity of $1 at Compound Interest
Year 6% 10% 12% 15% 20%
1 0.943 0.909 0.893 0.870 0.833
2 1.833 1.736 1.690 1.626 1.528
3 2.673 2.487 2.402 2.283 2.106
4 3.465 3.170 3.037 2.855 2.589
5 4.212 3.791 3.605 3.353 2.991
6 4.917 4.355 4.111 3.785 3.326
7 5.582 4.868 4.564 4.160 3.605
8 6.210 5.335 4.968 4.487 3.837
9 6.802 5.759 5.328 4.772 4.031
10 7.360 6.145 5.650 5.019 4.192

a. Determine the net present value for the two machines. Use the table of present values of an annuity of $1 above. Round to the nearest dollar.

Sewing Machine Packing Machine
Present value of annual net cash flows $ $
Amount to be invested $ $
Net present value $ $

b. Determine the present value index for the two machines. If required, round your answers to two decimal places.

Sewing Machine Packing Machine
Present value index

c. If Diamond and Turf has sufficient funds for only one of the machines and qualitative factors are equal between the two machines, in which machine should it invest? (If both present value indexes are the same, either machine will grade as correct.)
Sewing Machine
Answer #1
a
Sewing Machine Packing Machine
Present value of annual net cash flows 320587 151882
Amount to be invested 266100 132100
Net present value 54487 19782
b Sewing Machine Packing Machine
Present value index 1.20 1.15
c
Sewing Machine should be selected
Workings:
Annual cash flows Sewing Machine = (252-140)*0.42*1400= $65856
Annual cash flows Packing Machine = 26*1200= $31200
Sewing Machine Packing Machine
Annual cash flows 65856 31200
X PV factor @ 10% 4.868 4.868
Present value of annual net cash flows 320587 151882
Present value of annual net cash flows 320587 151882
Divided by Amount to be invested 266100 132100
Present value index 1.20 1.15